A thermodynamic system consists of an ideal gas in a cylinder. The gas is compressed from 5.68 L to 2.96 L at a constant pressure of 7.5 × 104 Pa. What work is involved in this process?
1
Expert's answer
2017-02-11T09:57:09-0500
A(work) = p*(V2 - V1) = 7.5*10^4 Pa*(2.96 L - 5.68 L)*10^-3 m^3/L = - 204 J